OpenStudy (anonymous):

GEOMETRY PROBLEM HELP PLEASE!! BE trisects angle ABD BF trisects angle DBC m

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910 can you help me?

jimthompson5910 (jim_thompson5910):

Let m<DBC = x If m<ABD : m<DBC = 4:1, then m<ABD = 4x

jimthompson5910 (jim_thompson5910):

since angles ABD and DBC are supplementary, we know that ABD + DBC = 180 4x + x = 180 x = ???

OpenStudy (anonymous):

x= 36

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so ABD = 4*36 = 144 and DBC = 36 degrees

jimthompson5910 (jim_thompson5910):

hopefully that's enough info to find the possible measures of EBF

OpenStudy (anonymous):

yea im still not sure?

jimthompson5910 (jim_thompson5910):

BE trisects angle ABD so what are the possible measures of ABE?

OpenStudy (anonymous):

72

OpenStudy (anonymous):

i dont know anything else?

jimthompson5910 (jim_thompson5910):

what was angle ABD again

OpenStudy (anonymous):

144

jimthompson5910 (jim_thompson5910):

you are trisecting that angle (ie cutting it into 3 equal pieces)

jimthompson5910 (jim_thompson5910):

144/3 = 48 degrees

jimthompson5910 (jim_thompson5910):

so each piece is 48 degrees that means that ABE is either 48 degrees or 48*2 = 96 degrees this is because we don't know if BE is closer to A or closer to D

jimthompson5910 (jim_thompson5910):

so again, angle ABE is either 48 degrees or 96 degrees

jimthompson5910 (jim_thompson5910):

I'll let you figure out the possible measures of EBD

OpenStudy (anonymous):

it can also be either 96 or 48

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now find the possible measures of DBF

OpenStudy (anonymous):

either be 24 or 12

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now add up the possible measures of EBD and the possible measures of DBF to get the possible measures of EBF

OpenStudy (anonymous):

72 60 120 108 are the choices

jimthompson5910 (jim_thompson5910):

good, you got them all